Which of these taxpayers has a casualty or theft gain? *

Aaron's personal-use vehicle (FMV $9,000, basis $15,000) was stolen and never recovered. His insurance reimbursed him $8,000 for the loss.

Elaine's main home (FMV $125,000, basis $105,000) was damaged by wind in a hurricane that was declared a federal disaster. Her insurance reimbursed her $12,000 for the damages.

Kirk's main home (FMV $250,000, basis $225,000) was damaged by a wildfire that was declared a federal disaster. His insurance reimbursed him $50,000 for the damages.

Gwen's diamond engagement ring (FMV $5,000, basis $2,500) was stolen and never recovered. Her insurance reimbursed her $4,000 for the loss.
Answer #1

If you receive an insurance payment or other reimbursement that is more than your adjusted basis in the destroyed, damaged, or stolen property, you have a gain from the casualty or theft. Your gain is figured as follows,

The amount you receive minus Your adjusted basis in the property at the time of the casualty or theft.

While you own the property, various events may take place that change your basis. Some events, such as additions or permanent improvements to the property, increase basis. Others, such as earlier casualty losses and depreciation deductions, decrease basis. When you add the increases to the basis and subtract the decreases from the basis, the result is your adjusted basis.

Since, the question doesn't specify anything, we assume that the basis given is Adjusted basis.

1. Aaron,

Amount received from insurance - Adjusted Basis of property = 8000-15000 = -7000$

2. Elaine,

Amount received from insurance - Adjusted Basis of property = 12000-105000 = -93000$

3. Kirk,

Amount received from insurance - Adjusted Basis of property = 50000-225000 = -175000$

4.Gwen,

Amount received from insurance - Adjusted Basis of property = 4000-2500 = 1500$ Gain

Therefore, the only person who has a casualty or theft gain from the above four cases is Gwen since the amount received from insurance is more than the adjusted basis of her diamond engagement ring. Her gain is 1500$.